A merchant wishes to mix peanuts costing Php 0.70 per gram with cashews costing Php 1.50 per gram obtaining 50 grams mixture which cost Php 1.20 per gram. How many grams of each type should be mixed?
Answer by addingup(3677)

Peanuts: 0.7
Cashews : 1.5
Mix: 1.2
Let the amount of peanuts be x. Then, the amount of cashews will be 50-x
:
0.7x+1.5(50-x) = 1.2(50)
0.7x+75-1.5x = 60
-0.8x = -15 Divide both sides by 0.8 and remember -/- = +
x = 18.75 grams of peanuts, and:
50-18.75 = 31.25 grams of cashews